OpenStudy (anonymous):

Find an equation of variation in which Y varies as X and the following is true Y=40 and X=38

OpenStudy (amistre64):

y = k*x

OpenStudy (amistre64):

40 = k * 38

OpenStudy (amistre64):

k = 40/38 = 20/19 y = (20/19)x